OpenStudy (anonymous):

An ancestor family tree shows that the farther you go back in time, the more ancestors you have. You have 16 great-great-grandparents in the 5th generation back, and you have about half a million ancestors who lived 500 years ago (the 20th generation back). The strange thing is, this makes it seem like the population of the world must be getting smaller, when we know it is actually getting bigger all the time. Which choice explains what is wrong with using an ancestor family tree to argue that the population of the world used to be bigger than it is now?

OpenStudy (anonymous):

these r the answer choices: A)Geometric sequences cannot be extended backward in time. B)The ratio of the geometric sequence is 1/2 , not 2. C)The first term of the geometric sequence cannot be determined. D)Many of the ancestors have died. E)The ancestor tree ignores the existence of brothers and sisters.

OpenStudy (amistre64):

A{1} = 2 ancestors A{2} = 2 + 2*2 ancestors A{3} = 2 + 2^2 + 2^3 ancestors A{4} = 2 + 2^2 + 2^3 + 2^4 ... we are doing in total right?

OpenStudy (anonymous):

no, i dont need this one solved i just need to know which of the answer choices to pick

OpenStudy (amistre64):

well; world population isnt a matter of who your ancestors are; its a matter of who the children are right? my best guess is d

OpenStudy (amistre64):

well; e or d aound good to me

OpenStudy (anonymous):

the answer was E but oh well

OpenStudy (amistre64):

i typoed it and said "d", but rationalized "e" ... :)

OpenStudy (anonymous):

also sibling share same ancestor

OpenStudy (anonymous):

That model only work if every parent at every generation have only one kid

OpenStudy (amistre64):

\begin{array}l\color{red}{\normalsize\text{t}}\color{orange}{\normalsize\text{h}}\color{#9c9a2e}{\normalsize\text{a}}\color{green}{\normalsize\text{t}}\color{blue}{\normalsize\text{s}}\color{purple}{\normalsize\text{ }}\color{purple}{\normalsize\text{n}}\color{red}{\normalsize\text{o}}\color{orange}{\normalsize\text{t}}\color{#9c9a2e}{\normalsize\text{ }}\color{green}{\normalsize\text{a}}\color{blue}{\normalsize\text{n}}\color{purple}{\normalsize\text{ }}\color{purple}{\normalsize\text{o}}\color{red}{\normalsize\text{p}}\color{orange}{\normalsize\text{t}}\color{#9c9a2e}{\normalsize\text{i}}\color{green}{\normalsize\text{o}}\color{blue}{\normalsize\text{n}}\color{purple}{\normalsize\text{ }}\color{purple}{\normalsize\text{!}}\color{red}{\normalsize\text{!}}\color{orange}{\normalsize\text{ }}\color{#9c9a2e}{\normalsize\text{ }}\color{green}{\normalsize\text{:}}\color{blue}{\normalsize\text{)}}\color{purple}{\normalsize\text{}}\end{array}

OpenStudy (anonymous):

E

OpenStudy (anonymous):

Here is graph I made modeling ancestry population if we assume that everyone have 3 offsprings(average). You will see that population grow each generation forward. https://docs.google.com/drawings/pub?id=1cFwCUuX0q8iej-o2m2v4ianZ9seVfTLWmMfdN0X_li8&w=960&h=720

OpenStudy (anonymous):

Question?

OpenStudy (anonymous):

not until my teacher resets my quiz my quiz